# Antiferromagnetism and spin excitations in a two-dimensional non-Hermitian Hatano-Nelson flux model

### v1

@@ III Dynamic spin susceptibility
To investigate the stability of the long-range $(\pi,\pi)$ AF order and the properties of the AF state we continue further by analyzing the longitudinal and transverse (spin wave) fluctuations of the magnetically ordered state. The bare dynamical spin susceptibility for the transverse, ++-+ -, and the longitudinal, $zz$, components is given by
[…]
where $lm=+-,zz$, with $S^{\pm}=S^{x}\pm iS^{y}$ corresponding to the usual spin ladder operators.
Since our non-Hermitian model consists of two sublattices from the very beginning, there are no extra umklapp terms appearing in the AF state. Therefore, within a random phase approximation (RPA) both components of the susceptibility can be expressed as,
@@
where $c=\pm$ are the band indices, $f(\varepsilon)$ is the Fermi-Dirac distribution and the prime refers to the sum over the (reduced) sBZ.
@@
and $\delta\bm{q}=\bm{q}-\bm{Q}_{\text{AF}}$. Equation (20) has poles at $\Omega(\bm{q})=\pm i\sqrt{c_{2}/c_{1}}\delta q_{x}\delta q_{y}$, which are purely imaginary, in contrast with ordinary forms of antiferromagnetic excitations featuring a real spectrum linear in the magnitude of the wave vector. On the other hand, an imaginary dispersion relation of the form $\Omega(q)=\pm c_{s}|\bm{q}|-iDq^{2}$ with $c_{s},D>0$ describes AF spin waves including dissipation from heat diffusion [27]. In our case the real part vanishes and the negative sign corresponds to a purely diffusive mode of a $d$-wave type. However, the positive sign leads to the opposite of a damping, so both signs taken together can be interpreted as being associated to gain and loss, in a behavior inherent to non-Hermitian systems. Usually, diffusive modes of the ordinary type naturally occur for the paramagnetic state [27]. The uncovered gain and loss process is finite only in a small interval of $W$ values and quickly approaches zero as $W$ increases. Furthermore, the coefficients $c_{1}$ and $c_{2}$ are well defined only in the region of unbroken $\mathcal{PT}$ symmetry ($W>t$). This means that the gain and loss mechanisms in this case take place even in the region where $\mathcal{PT}$ symmetry is unbroken.
The spin waves remain nearly flat also in the case of the semimetal and do not interfere with the particle-hole continuum, which is also gapless in this case. We conclude that the long-range $(\pi,\pi)$ AF order is highly frustrated due to nearly flat spin wave excitations spectrum. Note that the potential instability of the system towards the stripe-type order $(\pi,0)$ (or $(0,\pi)$) could be seen from the properties of the tight-binding dispersion at $\bm{Q}_{\text{X}}=(\pi,0)$, $\varepsilon_{\bm{k}+\bm{Q}_{\text{X}}}=i\varepsilon_{\bm{k}}$, which can be regarded as non-Hermitian nesting.
@@ IV Conclusions
To conclude, we have formulated the non-Hermitian flux model where each plaquette on a square lattice consists of a spinful, maximally asymmetric Hatano-Nelson model. As a consequence, the model is characterized by the clock- and anticlockwise asymmetric hoppings on a plaquette within the square lattice and resembles the flux phase discussed in the context of $U(1)$ spin liquid phases of AF $SU(N)$ Heisenberg models in the large $N$ limit [22]. At half-filling the non-interacting model describes a metal characterized by a Fermi surface with simultaneous exceptional lines separating the real- and complex-valued branches of the energy spectrum. After including an on-site Coulomb repulsion we analyzed the formation of a long-range AF phase with $(\pi,\pi)$ ordering wave vector at half-filling. We find that for the AF insulator and semimetal the AF order restores the $\mathcal{PT}$ symmetry with a real-valued energy spectrum. The transverse spin excitations (spin waves) are gapless and nearly flat along the M – Y and M – X directions of the BZ indicating potential frustration of $(\pi,\pi)$ AF order towards $(\pi,0)$ or $(0,\pi)$ states. This instability was demonstrated to be associated with the emergence of diffusive modes of a $d$-wave type. However, the diffusion is in this case does not only include a dissipative contribution (loss), but also an enhancement mode (gain), which is a consequence of the non-Hermiticity of the model. Moreover, this behavior occurs in the region of unbroken $\mathcal{PT}$ symmetry. Usually the non-standard behavior of non-Hermitian $\mathcal{PT}$ symmetric systems is always associated with the region where the $\mathcal{PT}$ symmetry is broken. Here we are dealing with a fundamentally new situation, where the phenomena of gain and loss occur even in the region of unbroken $\mathcal{PT}$ symmetry, as the evidenced by the behavior of the dynamic susceptibility. This indicates that even in the region of unbroken $\mathcal{PT}$ symmetry, one can expect phenomena related to the interaction of the system with the environment.

### v2

@@ III Dynamic spin susceptibility
To investigate the stability of the long-range $(\pi,\pi)$ AF order and the properties of the AF state we continue further by analyzing the longitudinal and transverse (spin wave) fluctuations of the magnetically ordered state. Usually, models with non-Hermitian excitation require a modified response function [35, 36, 37, 38, 39]. However, in our case, although the model itself is non-Hermitian, the excitation is Hermitian. Therefore, we can use the usual definition for bare dynamical spin susceptibility for the transverse, ++-, and the longitudinal, $zz$, components
$\chi^{lm}_{0}(\bm{q},\bm{q}^{\prime},\Omega)=\int dte^{i\Omega t}\langle T_{t}S^{l}_{\bm{q}}(t)S^{m}_{-\bm{q}^{\prime}}(0)\rangle,$ (17)
where lm=+,zzlm=+-,$z$, with $S^{\pm}=S^{x}\pm iS^{y}$ corresponding to the usual spin ladder operators, $\Omega$ is a frequency and $T_{t}$ is a time ordering operator.
Within the random phase approximation (RPA) framework, the total susceptibility is expressed as a sum of bubble diagrams, where each bubble represents the susceptibility of a non-interacting system calculated with respect to the mean-field ground state.
Normally RPA summation is a $2\times 2$ matrix problem since the antiferromagnetic ordering at $\bm{q}=(\pi,\pi)$ doubles the unit cell, but our non-Hermitian model consists of two sublattices from the very beginning. As a consequence there are no extra umklapp terms appearing in the AF state. Therefore, within RPA both components of the susceptibility can be expressed as,
@@
where $c=\pm$ are the band indices, $f(\varepsilon)$ is the Fermi-Dirac distribution and $V$ is the volume. The prime refers to the sum over the (reduced) sBZ. For details, see Appendix A.
@@
In Appendix B we show that the same result can be obtained using the partition function in bi-orthogonal basis.
@@
and $\delta\bm{q}=\bm{q}-\bm{Q}_{\text{AF}}$. Equation (23) has poles at $\Omega(\bm{q})=\pm i\sqrt{c_{2}/c_{1}}\delta q_{x}\delta q_{y}$, which are purely imaginary, in contrast with ordinary forms of antiferromagnetic excitations featuring a real spectrum linear in the magnitude of the wave vector. On the other hand, an imaginary dispersion relation of the form $\Omega(q)=\pm c_{s}|\bm{q}|-iDq^{2}$ with $c_{s},D>0$ describes AF spin waves including dissipation from heat diffusion [41]. In our case the real part vanishes and the negative sign corresponds to a purely diffusive mode of a $d$-wave type. However, the positive sign leads to the opposite of a damping, so both signs taken together can be interpreted as being associated to gain and loss, in a behavior inherent to non-Hermitian systems. Usually, diffusive modes of the ordinary type naturally occur for the paramagnetic state [41]. The uncovered gain and loss process is finite only in a small interval of $W$ values and quickly approaches zero as $W$ increases. Furthermore, the coefficients $c_{1}$ and $c_{2}$ are well defined only in the region of unbroken $\mathcal{PT}$ symmetry ($W>t$), in which case the energy spectrum (10) is real. Usually in this case we are not dealing with such non-unitary phenomena as gain and loss. However, for our model the region of unbroken $\mathcal{PT}$ symmetry corresponds to the imaginary dispersion relation for spin waves. This means that the gain and loss mechanisms in this case take place even in the region where $\mathcal{PT}$ symmetry is unbroken.
The spin waves remain nearly flat also in the case of the AF semimetal and do not interfere with the particle-hole continuum, which is also gapless in this case. We conclude that the long-range $(\pi,\pi)$ AF order is highly frustrated due to nearly flat spin wave excitations spectrum. Note that the potential instability of the system towards the stripe-type order $(\pi,0)$ (or $(0,\pi)$) could be seen from the properties of the tight-binding dispersion at $\bm{Q}_{\text{X}}=(\pi,0)$, $\varepsilon_{\bm{k}+\bm{Q}_{\text{X}}}=i\varepsilon_{\bm{k}}$, which can be regarded as non-Hermitian nesting.
@@ IV Conclusions
To conclude, we have formulated the non-Hermitian flux model where each plaquette on a square lattice consists of a spinful, maximally asymmetric Hatano-Nelson model. As a consequence, the model is characterized by the clock- and anticlockwise asymmetric hoppings on a plaquette within the square lattice and resembles the flux phase discussed in the context of $U(1)$ spin liquid phases of AF $SU(N)$ Heisenberg models in the large $N$ limit [29]. At half-filling the non-interacting model describes a metal characterized by a Fermi surface with simultaneous exceptional lines separating the real- and complex-valued branches of the energy spectrum. After including an on-site Coulomb repulsion we analyzed the formation of a long-range AF phase with $(\pi,\pi)$ ordering wave vector at half-filling. We find that for the AF insulator the AF order restores the $\mathcal{PT}$ symmetry with a real-valued energy spectrum. At the same time, for the AF metal, $\mathcal{PT}$ symmetry can be broken for certain values of momentum.
The transverse spin excitations (spin waves) are gapless and nearly flat along the M – Y and M – X directions of the BZ indicating potential frustration of $(\pi,\pi)$ AF order towards $(\pi,0)$ or $(0,\pi)$ states. This instability was demonstrated to be associated with the emergence of diffusive modes of a $d$-wave type. However, the diffusion is in this case does not only include a dissipative contribution (loss), but also an enhancement mode (gain), which is a consequence of the non-Hermiticity of the model. Moreover, this behavior occurs in the region of unbroken $\mathcal{PT}$ symmetry. Usually the non-standard behavior of non-Hermitian $\mathcal{PT}$ symmetric systems is always associated with the region where the $\mathcal{PT}$ symmetry is broken. Here we are dealing with a fundamentally new situation, where the phenomena of gain and loss occur even in the region of unbroken $\mathcal{PT}$ symmetry, as the evidenced by the behavior of the dynamic susceptibility. This indicates that even in the region of unbroken $\mathcal{PT}$ symmetry, one can expect phenomena related to the interaction of the system with the environment.
@@
We believe that there are two fairly simple ways to implement our model. The first method consists in the experimental realization of a non-Hermitian Hamiltonian featuring a non-Hermitian skin effect (NHSE) in a momentum lattice using ultracold atoms [42, 43, 44, 45]. Namely, there is a possibility of creating a chain of coupled Aharonov-Bohm (AB) rings along a synthetic momentum lattice [9] using Bose-Einstein condensate of 87Rb atoms [46, 47, 48].
The second method involves ready-made two-dimensional non-reciprocal topological electric circuits [14, 13, 15] with slightly modified hopping according to the definition of the flux model.
@@ Acknowledgements.
Appendix A Derivation of the dynamical spin susceptibility
To calculate the bare dynamical spin susceptibility for the transverse and the longitudinal components, we first need to obtain the matrix elements of the Green’s function for the mean-field Hamiltonian (12)
$\hat{G}(\bm{k},i\omega_{n})$ $=$ $(-i\omega_{n}\mathbbm{1}-\hat{H}_{\text{MF}}(\bm{k}))^{-1},$ (26)
in order to obtain matrix elements of Green’s function,
$G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},i\omega_{n})=-\frac{u_{\bm{k},\sigma}}{i\omega_{n}-E^{+}_{\bm{k}}}-\frac{v_{\bm{k},\sigma}}{i\omega_{n}-E^{-}_{\bm{k}}},$ (27a)
$G_{b^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},i\omega_{n})=-\frac{v_{\bm{k},\sigma}}{i\omega_{n}-E^{+}_{\bm{k}}}-\frac{u_{\bm{k},\sigma}}{i\omega_{n}-E^{-}_{\bm{k}}},$ (27b)
$G_{a^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},i\omega_{n})=w_{\bm{k}}\left(\frac{1}{i\omega_{n}-E^{+}_{\bm{k}}}-\frac{1}{i\omega_{n}-E^{-}_{\bm{k}}}\right),$ (27c)
$G_{b^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},i\omega_{n})=x_{\bm{k}}\left(\frac{1}{i\omega_{n}-E^{+}_{\bm{k}}}-\frac{1}{i\omega_{n}-E^{-}_{\bm{k}}}\right),$ (27d)
where
$u_{\bm{k},\sigma}$ $=$ $\frac{1}{2}\left(1+\frac{\sigma W}{\sqrt{\varepsilon^{2}_{\bm{k}}+W^{2}}}\right),\>w_{\bm{k}}=\frac{t\cos k_{y}}{2\sqrt{\varepsilon^{2}_{\bm{k}}+W^{2}}},\hskip 14.22636pt$ (28)
$v_{\bm{k},\sigma}$ $=$ $\frac{1}{2}\left(1-\frac{\sigma W}{\sqrt{\varepsilon^{2}_{\bm{k}}+W^{2}}}\right),\>x_{\bm{k}}=\frac{t\cos k_{x}}{2\sqrt{\varepsilon^{2}_{\bm{k}}+W^{2}}}.$ (29)
We can now proceed directly to the definition of bare dynamical spin susceptibility (17) and write down the relation between the fermionic and spin operators,
$S^{+}_{\bm{q}}=\frac{1}{\sqrt{V}}\sum_{\bm{k}}{}^{{}^{\prime}}\left(a^{\dagger}_{\bm{k}+\bm{q},\uparrow}a_{\bm{k},\downarrow}+b^{\dagger}_{\bm{k}+\bm{q},\uparrow}b_{\bm{k},\downarrow}\right),$ (30a)
$S^{-}_{\bm{q}}=\frac{1}{\sqrt{V}}\sum_{\bm{k}}{}^{{}^{\prime}}\left(a^{\dagger}_{\bm{k}+\bm{q},\downarrow}a_{\bm{k},\uparrow}+b^{\dagger}_{\bm{k}+\bm{q},\downarrow}b_{\bm{k},\uparrow}\right),$ (30b)
$S^{z}_{\bm{q}}=\frac{1}{\sqrt{V}}\sum_{\bm{k}}{}^{{}^{\prime}}\Big{(}a^{\dagger}_{\bm{k}+\bm{q},\uparrow}a_{\bm{k},\uparrow}-a^{\dagger}_{\bm{k}+\bm{q},\downarrow}a_{\bm{k},\downarrow}$
$\hskip 14.22636pt+b^{\dagger}_{\bm{k}+\bm{q},\uparrow}b_{\bm{k},\uparrow}-b^{\dagger}_{\bm{k}+\bm{q},\downarrow}b_{\bm{k},\downarrow}\Big{)}.$ (30c)
By substituting Eqs. (30) into the dynamical spin susceptibilities and using Vick’s theorem, we can represent Eq. (17) in terms of the corresponding matrix elements of the Green’s function,
$\chi^{+-}_{0}(\bm{q},\bm{q},i\Omega_{n})$ $=$ $\frac{1}{\beta V}\sum_{\bm{k},m}{}^{\prime}\bigg{[}G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},i\omega_{m})G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})+G_{b^{\dagger}_{\downarrow}b_{\downarrow}}(\bm{k},i\omega_{m})G_{b^{\dagger}_{\uparrow}b_{\uparrow}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})$ (31)
$+$ $G_{a^{\dagger}_{\downarrow}b_{\downarrow}}(\bm{k},i\omega_{m})G_{b^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})+G_{b^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},i\omega_{m})G_{a^{\dagger}_{\uparrow}b_{\uparrow}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})\bigg{]}$
and
$\chi^{zz}_{0}(\bm{q},\bm{q},i\Omega_{n})$ $=$ $\frac{1}{\beta V}\sum_{\bm{k},m,\sigma}{}^{\prime}\bigg{[}G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},i\omega_{m})G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})+G_{b^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},i\omega_{m})G_{b^{\dagger}_{\sigma}b_{\sigma}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})$ (32)
$+$ $G_{a^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},i\omega_{m})G_{b^{\dagger}_{\sigma}a_{\sigma}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})+G_{b^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},i\omega_{m})G_{a^{\dagger}_{\sigma}b_{\sigma}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})\bigg{]},$
where $\Omega_{n}=2n\pi/\beta$ and $\omega_{n}=(2n+1)\pi/\beta$, $n\in\mathbb{Z}$ are bosonic and fermionic Matsubara frequencies, respectively and $\beta=1/T$ is the inverse of temperature. As a result, we have a set of bubble diagrams. Let’s make detailed calculations of one of them. The rest of the diagrams can be calculated in a similar way.
Figure 7: The integration contour $C$ is represented by a blue line. The dashed lines indicate two branch cuts in the complex plane: $\text{Im}(z)=0$ and $\text{Im}(z+i\Omega_{n})=0$. The red crosses correspond to the poles of Fermi-Dirac distribution function.
First we can replace the Matsubara frequency summation by a contour integral
$\frac{1}{\beta}\sum_{m}G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},i\omega_{m})G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},i\omega_{m}+i\Omega_{n})$
$\hskip-14.22636pt=-\oint\limits_{C}\frac{dz}{2\pi i}f(z)G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},z)G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},z+i\Omega_{n}),$ (33)
where $f(z)$ is the Fermi-Dirac distribution function. The contour is shown in Fig. 7. In his contour integral, we need to avoid two horizontal branch cuts defined by $\text{Im}(z)=0$ and $\text{Im}(z+i\Omega_{n})=0$. To calculate the integral, we note that the semicircular arcs do not contribute to the integral, and therefore only the contribution from the horizontal lines remains. Now we can parameterize the upper branch cut as $z=\epsilon+i0^{+}$, $z=\epsilon-i0^{+}$ and the lower cut as $z=\epsilon-i\Omega_{n}+i0^{+}$, $z=\epsilon-i\Omega_{n}-i0^{+}$. Thus we have
$-\oint\limits_{C}$ $\frac{dz}{2\pi i}f(z)G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},z)G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},z+i\Omega_{n})$
$=\int\limits_{-\infty}^{\infty}\frac{d\epsilon}{2\pi i}f(\epsilon)G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},\epsilon+i\Omega_{n})$
$\times\left[G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},\epsilon+i0^{+})-G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},\epsilon-i0^{+})\right]$
$+\int\limits_{-\infty}^{\infty}\frac{d\epsilon}{2\pi i}f(\epsilon)G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},\epsilon-i\Omega_{n})$
$\times\left[G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},\epsilon+i0^{+})-G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},\epsilon-i0^{+})\right].$ (34)
Now we can move on to retarded functions by making the substitution: $i\Omega_{n}=\Omega+i0^{+}$ and use the following relation for Green’s functions
$G_{a^{\dagger}_{\sigma}a_{\sigma}}$ $(\bm{k},\epsilon+i0^{+})-G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},\epsilon-i0^{+})$
$=2i\text{Im}[G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},\epsilon+i0^{+})].$ (35)
In turn, for the imaginary parts of the Green’s functions (27) we have
Im $[G_{a^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},\omega+i0^{+})]$
$=\pi\left[u_{\bm{k},\sigma}\delta(\omega-E^{+}_{\bm{k}})+v_{\bm{k},\sigma}\delta(\omega-E^{-}_{\bm{k}})\right],$ (36a)
Im $[G_{b^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},\omega+i0^{+})]$
$=\pi\left[v_{\bm{k},\sigma}\delta(\omega-E^{+}_{\bm{k}})+u_{\bm{k},\sigma}\delta(\omega-E^{-}_{\bm{k}})\right],$ (36b)
Im $[G_{a^{\dagger}_{\sigma}b_{\sigma}}(\bm{k},\omega+i0^{+})]$
[…]
Im $[G_{b^{\dagger}_{\sigma}a_{\sigma}}(\bm{k},\omega+i0^{+})]$
[…]
Finally we can substitute Eqs. (27) and (36) into Eq. (34) and easily integrate expressions with delta functions. After that we obtain
$-\oint\limits_{C}$ $\frac{dz}{2\pi i}f(z)G_{a^{\dagger}_{\downarrow}a_{\downarrow}}(\bm{k},z)G_{a^{\dagger}_{\uparrow}a_{\uparrow}}(\bm{k}+\bm{q},z+i\Omega_{n})$
$=u_{\bm{k+q},\uparrow}u_{\bm{k},\downarrow}\frac{f(E^{+}_{\bm{k}})-f(E^{+}_{\bm{k+q}})}{\Omega+i0^{+}-E^{+}_{\bm{k+q}}+E^{+}_{\bm{k}}}$
$+v_{\bm{k+q},\uparrow}v_{\bm{k},\downarrow}\frac{f(E^{-}_{\bm{k}})-f(E^{-}_{\bm{k+q}})}{\Omega+i0^{+}-E^{-}_{\bm{k+q}}+E^{-}_{\bm{k}}}$
$+u_{\bm{k+q},\uparrow}v_{\bm{k},\downarrow}\frac{f(E^{-}_{\bm{k}})-f(E^{+}_{\bm{k+q}})}{\Omega+i0^{+}-E^{+}_{\bm{k+q}}+E^{-}_{\bm{k}}}$
$+v_{\bm{k+q},\uparrow}u_{\bm{k},\downarrow}\frac{f(E^{+}_{\bm{k}})-f(E^{-}_{\bm{k+q}})}{\Omega+i0^{+}-E^{-}_{\bm{k+q}}+E^{+}_{\bm{k}}}$. (37)
Performing similar calculations for the remaining bubble diagrams in Eqs. (31) and (32), after straightforward simplification using expressions (28) and (29), we obtain Eqs. (19) and (20).
Appendix B Partition function in the bi-orthogonal basis
In this appendix, we show that using bi-orthogonal basis leads to to the same mean-field equation as previous calculations using the Green’s function.
Although the matrix Hamiltonian (12) is non-Hermitian, it is pseudo-Hermitian,
$\hat{\eta}(\bm{k})\hat{H}_{\text{MF}}(\bm{k})=\hat{H}^{\dagger}_{\text{MF}}(\bm{k})\hat{\eta}(\bm{k}),$ (38)
$\hat{\eta}(\bm{k})=\begin{bmatrix}\cos k_{x}/\cos k_{y}&0\\ 0&1\end{bmatrix}$. (39)
The bi-orthogonal basis includes the following set of right and left eigenvectors
[…]
and
$N^{\pm}_{\bm{k},\sigma}=1/\sqrt{1+\left(\mp\sigma W+\sqrt{\varepsilon^{2}_{\bm{k}}+W^{2}}\right)/\varepsilon^{2}_{\bm{k}}}.$ (41)
One can verify that
[…]
and
[…]
where $\alpha,\beta=\pm$. Calculating the partition function in terms of vectors (40a) and (40b) leads to the standard expression,
$Z=e^{-\beta V\frac{W^{2}}{2U}}\prod_{\bm{k}}{}^{{}^{\prime}}\left(1+e^{-\beta E^{+}_{\bm{k}}}\right)\left(1+e^{-\beta E^{-}_{\bm{k}}}\right),$ (44)
which we use to obtain the free energy density,
[…]
Now from the condition of free energy density minimization $\partial f/\partial W=0$ we obtain self-consistent equation that completely coincides with Eq. (22).